5. 20 pts Prove that if n is an odd positive integer, then n 1 (mod 8)

this needs to be proven using a direct proof, proof by contraposition, or a proof by contradiction. YOU MUST SHOW ALL STEPS LABEL THEM AND SPACE THEM WELL PLEASE. ALSO INCLUDE ALL DEFINITIONS USED . FORMAL PROOFS ONLY PLEASE
Answer #1

Let i be an integer number

Even Integer number = 2i

Odd Integer Number = 2i + 1

where i=1,2,3,\dots

Odd Integer Number n=2i+1

n^2=(2i+1)^2 \newline n^2 = (2i+1)^2 = 4i^2 +4i +1 \newline (n^2 - 1) = 4i(i +1)

If i is an even number then i(i+1) is also even value

If i is an odd number then i(i+1) is also even value

So here l\epsilon Z such that i(i+1) = 2l

n^2-1 = 4 i (i+1) = 4\times 2l =8l

n^2=8l+1

Taking Mod 8 on Both Sides

n^2(mod\hspace{1em} 8)=(8l+1)(mod \hspace{1em} 8) \newline =(8l)(mod \hspace{1em} 8) + 1(mod \hspace{1em} 8) \newline =0 (mod \hspace{1em} 8) + 1(mod \hspace{1em} 8) \newline \equiv 1(mod \hspace{1em} 8)